A square sign must have a perimeter of no more than 225 inches. what lengths are acceptable for the sides of the sign?
Answer by fcabanski(1391)

Check my profile and contact me for one on one online tutoring.


The formula for the perimeter (P) of a square with sides length S is:

P = 4S. Divide both sides by 4 to find the length of a side.


P/4 = S.


In this problem we're given a maximum P (225 inches). So S must be no bigger than a length that will result in 225 inches. That is:

S <= P/4 Which is S <= 225/4. That gives: S <= 56.25 inches.


Sides can be any length equal to 56.25 inches, or smaller than 56.25 inches.